OpenStudy (anonymous):

what is the curve of intersection of the surfaces x^2+z^2 = 1; and y = x^2

OpenStudy (anonymous):

How would you come up with a parametrization of this curve?

OpenStudy (anonymous):

or let's try it this way. the curve of intersection of surfaces x^2+z^2 = 1; and y = x^2 can be parametrized : r = (cost, cos(t)^2, sint) ..........(vector).. how did this happen?

OpenStudy (anonymous):

Ah, I figured it out in case anyone was interested, wasn't really that hard. but i hate parametrizations. x^2=1-z^2 x^2 = y let x = cost then. y = cos^(2)t and z = sqrt(1-(cost)^2)) the function on the inside of the square root is alternately sin(t)^2 so z = sint so r(t) = (cost, cos(t)^2, sint)